OpenStudy (anonymous):

The bottle of vanilla Diann uses in her baking contains 0.473 liters. If her cookie recipe calls for 1 teaspoon of vanilla ( 1 teaspoon = 5 mililiters ), how many batched of cookies can Diann make with one bottle of vanilla? How many mililliters will Diann have left?

OpenStudy (anonymous):

(x liters / 1 bottle) (1000 milliliters / 1 liter) (1 teaspoon / 5 milliliters) (1 batch / 1 teaspoon). You will get a decimal answer for the # of batches that you will have to truncate.

OpenStudy (anonymous):

What am i supposed to do?

OpenStudy (anonymous):

First, notice how units cancel in the denominator and numerator. This is basically a long string of multiplication, a lot of the factors are "multiplying by one" to preserve the identity since multiplying by one preserves the answer. The first thingto do is to understand this explanation. The second is to try to set it up on your own without looking. Then, the thrird thingto do is substitute for x and mulitply it out, cancelling units .

OpenStudy (anonymous):

im sorry but I am confused...

OpenStudy (anonymous):

(a units of m / b units of n) (c units of n / d units of p) = (ac units of m) / (bd units of p). Notice how "units of n" was in both the numerator and denominator, so it cancels out. This is exactly what is going on in the equation I set up for you. Study this simpler example and apply it to the original problem. Any more help, and I would just be giving you an answer.

OpenStudy (anonymous):

i am not understanding the whole unit thing.

OpenStudy (anonymous):

(Unit of something) / (Unit of something) = 1. You can multiply anything by 1 and the "anything" stays the same. (one inch) / (one inch) = 1. When you have the same thing in the numerator and the denominator, it is "1" which is the multiplicative identity, which means that the original thing being multiplied stays the same.

OpenStudy (anonymous):

I am really sorry but this is so confusing...

OpenStudy (anonymous):

I can't explain it any better or in any more detail without giving an answer. Perhaps a live tutor can help where the tutor can see your work in progress. If you can attempt a start here and write something down after thinking about it, perhaps I can try to help further. Otherwise, I don't know how to help you further at this point.

OpenStudy (anonymous):

Thank you so much for your help. I will try my hardest and see what I can get... Thanks. :)